# Direct and Converse Theorems in Estimating Signals with Sublinear Sparsity

### v1

@@ Abstract
This paper addresses the estimation of signals with sublinear sparsity sent over the additive white Gaussian noise channel. This fundamental problem arises in designing denoisers used in message-passing algorithms for sublinear sparsity. The main results are direct and converse theorems in the sublinear sparsity limit, where the signal sparsity grows sublinearly in the signal dimension as the signal dimension tends to infinity. As a direct theorem, the maximum likelihood estimator is proved to achieve vanishing square error in the sublinear sparsity limit if the noise variance is smaller than a threshold. As a converse theorem, all estimators cannot achieve square errors smaller than the signal power under a mild condition if the noise variance is larger than another threshold. In particular, the two thresholds coincide with each other when non-zero signals have constant amplitude. These results imply the asymptotic optimality of an existing separable Bayesian estimator used in approximate message-passing for sublinear sparsity.
@@ I-A Motivation
with noise variance $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$ for some $\sigma^{2}>0$. The signal support $\mathcal{S}_{\boldsymbol{x}}=\{n\in\mathcal{N}=\{1,\ldots,N\}:x_{n}\neq 0\}$ is assumed to be sampled from all possible sets $\{\mathcal{S}\subset\mathcal{N}:|\mathcal{S}|=k\}$ of size $k$ uniformly and randomly. In particular, this paper focuses on signals with sublinear sparsity: The signal sparsity kk—number of non-zero elements in $\boldsymbol{x}$—tends to infinity sublinearly in $N$ as $N\to\infty$. The goal is estimation of the sparse signal vector $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$. This simple problem arises in designing denoisers used in message-passing algorithms for sublinear sparsity [1, 2].
The noise power scaling $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$ in (1) is a key point in estimation of the sparse signal vector. Conventional research for sublinear sparsity [3, 4, 5, 6] postulated $N^{-1}\mathbb{E}[\|\boldsymbol{\omega}\|_{2}^{2}]={\cal O}(1)$. This normalization implies that error-free estimation is impossible as long as the non-zero signal power is ${\cal O}(1)$. To evaluate the performance of an estimator $\hat{\boldsymbol{x}}\in\mathbb{R}^{N}$, the normalized square error $N^{-1}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}$ was used. For the scaling in this paper, on the other hand, $k^{-1}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}$ normalized by sparsity kk—simply called square error—should be considered [1, 2].
@@
When the separable Bayesian estimator is used as a denoiser, both approximate message-passing (AMP) [1] and orthogonal AMP (OAMP) [2] for sublinear sparsity require strictly larger sample complexity than information-theoretically optimal complexity [7, 8, 9, 10, 11, 12]. This suboptimality is in contrast to the optimality of AMP and OAMP for conventional linear sparsity $k/N\to\rho$ for some $\rho\in(0,1]$: For linear sparsity, AMP [13, 14] was proved to be Bayes-optimal [15, 16, 17] via state evolution [18, 19, 20, 21] when sparse signals are compressed with zero-mean independent and identically distributed (i.i.d.) sub-Gaussian matrices. Similarly, OAMP [22] or equivalently vector AMP (VAMP) [23] was proved to be Bayes-optimal [24, 25] for right-orthogonally invariant matrices via state evolution [23, 26]. See [27, 28] for a generalization of right-orthogonally invariant matrices. In contrast to these optimality for linear sparsity, both AMP and OAMP are suboptimal for sublinear sparsity [1, 2].
@@
  • In the direct part, i.e. achievability of vanishing square error, Gallager’s bound [29] is available for evaluating the performance of the maximum likelihood (ML) estimator. In contrast to a long heuristic proof for the separable Bayesian estimator [1], the proof of the direct part is simplified for the ML estimator.
@@
The converse part, i.e. impossibility of vanishing square error is proved via a hypothesis testing [34, 35]. In the hypothesis testing, we test whether the null hypothesis $\boldsymbol{y}\sim q(\boldsymbol{y})$ holds for some probability density function (pdf) $q(\boldsymbol{y})$ or the alternative hypothesis $\boldsymbol{y}\sim p(\boldsymbol{y})$. When the true marginal pdf $p(\boldsymbol{y})$ of $\boldsymbol{y}$ in (1) is indistinguishable from $q(\boldsymbol{y})$ independent of the channel model (1), it is intuitively impossible to estimate the signal vector $\boldsymbol{x}$. Inspired by [12], this paper utilizes a general relationship [36] between mutual information and square error to derive a lower bound on the square error, which depends on the Kullback-Leibler (KL) divergence $D(p\|q)$ of $q(\boldsymbol{y})$ from the truth $p(\boldsymbol{y})$. By proving that $D(p\|q)$ converges to zero, this paper proves a converse result.
@@ I-B Contributions
To present theoretical results of this paper, additional assumptions are summarized. Throughout this paper, non-zero elements of the signal vector are sampled from the set of $M$ discrete non-zero points $\mathcal{U}=\{u_{m}\in\mathbb{R}\setminus\{0\}:m\in\mathcal{M}=\{1,\ldots,M\}\}$ uniformly and randomly. Let $u_{\mathrm{min}}=\min_{m\in\mathcal{M}}|u_{m}|$ and $u_{\mathrm{max}}=\max_{m\in\mathcal{M}}|u_{m}|$. The set of all possible $k$-sparse signal vectors is written as $\mathcal{X}_{k}^{N}(\mathcal{U})=\{\boldsymbol{x}\in\mathbb{R}^{N}:|\mathcal{S}_{\boldsymbol{x}}|=k,x_{n}\in\mathcal{U}~\hbox{for all $n\notin\mathcal{S}_{\boldsymbol{x}}$}\}$.
@@
Let $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ denote the ML estimator of $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$ in (1). If $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ holds, then the square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]$ converges to zero in the sublinear sparsity limit: $k$ and $N$ tend to infinity with $\log k/\log N\to\gamma$ for some $\gamma\in[0,1)$.
Theorem 1 implies that the ML estimator is comparable to the separable Bayesian estimator [1] in the low noise regime $\sigma^{2}<u_{\mathrm{min}}^{2}/2$. Since the ML estimator can be suboptimal in terms of the square error, Theorem 1 makes no claims for high noise regime. This paper presents the following converse result for high noise regime:
@@
$\gamma<\min\left\{\frac{C_{0}}{C_{0}+\sigma^{2}/(u_{\mathrm{max}}u_{\mathrm{min}})},\frac{C_{1}}{C_{1}+3},\;\frac{1}{2}\right\},$ (3)
@@
To obtain a tight result, assume constant non-zero signals, i.e. $u_{\mathrm{min}}=u_{\mathrm{max}}$. The upper bound in (3) is strictly positive as long as $\sigma^{2}>u_{\mathrm{min}}^{2}/2$ is satisfied. Combining Theorems 1 and 2, this paper concludes that, for small $\gamma$, the square error converges to zero in the sublinear sparsity limit if and only if $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ holds. In other words, the separable Bayesian estimator [1] is asymptotically optimal for small $\gamma$.
The last contribution is numerical results. A non-separable Bayesian estimator is proposed via an approximation of the posterior mean estimator in the noiseless limit. The proposed estimator is numerically shown to outperform the ML estimator for all signal-to-noise ratios (SNRs) in terms of the square error. Furthermore, it is superior to the separable Bayesian estimator [1] in the high SNR regime. Unfortunately, the non-separable Bayesian estimator cannot improve the performance of AMP [1] for sublinear sparsity compared to the separable Bayesian estimator. The last result implies that AMP for sublinear sparsity has a performance bottleneck in another part rather than the type of denoisers.
@@ I-D Notation
The notation $p(X)$ denotes the pdf of an absolutely continuous random variable $X$ while $\mathbb{P}(\cdot)$ represents a probability measure. The Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ is denoted by $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. The function $Q(x)$ represents the complementary cumulative distribution function of the standard Gaussian distribution.
@@ II-A Maximum Likelihood Estimation
which minimizes the error probability $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{MAP}}\neq\boldsymbol{x})$. In terms of AMP [1] and OAMP [2] for sublinear sparsity, the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{PME}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ should be used to minimize the mean square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{PME}}\|_{2}^{2}]$. Nonetheless, this paper focuses on the ML estimator (6) because the optimality of the ML estimator is partially proved in the converse part.
The ML estimator (6) can be represented via a partition of indices $\mathcal{N}=\{1,\ldots,N\}$: For some $k_{*}\in\{0,\ldots,k\}$ decompose $\mathcal{N}$ into three disjoint sets $\mathcal{N}_{1}\subset\mathcal{N}$, $\mathcal{N}_{2}\subset\mathcal{N}\setminus\mathcal{N}_{1}$, and $\mathcal{N}_{3}=\mathcal{N}\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$ with cardinality $_{*}, kk$$k_{*}$, and NkN-k, respectively. The set $\mathcal{N}_{1}$ consists of indices that correspond to the top $k_{*}$ maximum received signals while $\mathcal{N}_{2}$ is composed of indices for the top $(k-k_{*})$ minimum received signals. The set $\mathcal{N}_{3}$ consists of the remaining indices.
@@
Let $\boldsymbol{y}]_{(n)}$ denote the $n$th maximum element of $\boldsymbol{y}$. To determine $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, define
@@ II-B Proof of Theorem 1
We consider $\mathcal{E}_{d}=\{\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\geq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}d\}$ with $d=\lceil k/\sqrt{\log(N/k)}\rceil$. Since $\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}$ is bounded from above by $\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}k$, we use the representation $\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]=\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}|\mathcal{E}_{d}]\mathbb{P}(\mathcal{E}_{d})+\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}|\mathcal{E}_{d}^{\mathsf{c}}]\mathbb{P}(\mathcal{E}_{d}^{\mathsf{c}})$ to obtain the following upper bound:
@@
The main challenge is evaluation of $\mathbb{P}(\mathcal{E}_{d})$. To evaluate this probability, we classify the signal space into types. Let $\boldsymbol{u}_{\boldsymbol{m}}=[u_{m_{1}},\ldots,u_{m_{k}}]^{\mathrm{T}}\in\mathcal{U}^{k}$ denote the vector of non-zero signal elements, with an index vector $\boldsymbol{m}\in\mathcal{M}^{k}$. For $\mathcal{S}\subset\mathcal{N}$, $\boldsymbol{m}\in\mathcal{M}^{|\mathcal{S}|}$, and $w\in\{0,\ldots,|\mathcal{S}|\}$, define the type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ of $|\mathcal{S}|$-sparse vectors in $\mathcal{X}_{|\mathcal{S}|}^{N}(\mathcal{U})$ as
@@
The type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ in (12) contains all possible $k$-sparse vectors that have non-zero elements $\boldsymbol{u}_{\boldsymbol{m}}$ and $w$ incorrectly detected positions for the signal vector $\boldsymbol{x}$. By definition, we have the following partition:
@@
We reduce the evaluation of $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$ to that of $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$ for $w^{\prime}\in\{0,\ldots,k-w\}$. The integer $w^{\prime}\in\{0,\ldots,k-w\}$ in $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ represents the number of incorrectly detected non-zero elements in $n\in\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}$. Thus, this joint event implies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\leq 2u_{\mathrm{max}}^{2}w+d_{\mathrm{max}}^{2}w^{\prime}\leq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}(w+w^{\prime})$. By definition, $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$ is affected only by the events satisfying $w+w^{\prime}\geq d$ for all $w\in\{0,\ldots,k\}$ and $w^{\prime}\in\{0,\ldots,k-w\}$:

### v3

@@ Abstract
This paper addresses the estimation of signals with sublinear sparsity sent over the additive white Gaussian noise channel. This fundamental problem arises in designing denoisers used in message-passing algorithms for sublinear sparsity. From a theoretical perspective, the main results are direct and converse theorems in the sublinear sparsity limit, where the signal sparsity grows sublinearly in the signal dimension as the signal dimension tends to infinity. As a direct theorem, the maximum likelihood estimator is proved to achieve vanishing square error in the sublinear sparsity limit if the noise variance is smaller than a threshold. This threshold is known to be achievable by an existing separable Bayesian estimator. As a converse theorem, all estimators cannot achieve square errors smaller than the signal power under a mild condition if the noise variance is larger than another threshold. In particular, the two thresholds coincide with each other when non-zero signals have constant amplitude. These results imply the asymptotic optimality of the existing separable Bayesian estimator used in approximate message-passing for sublinear sparsity. From a numerical perspective, a non-separable estimator is proposed via a heuristic approximation of the true posterior mean estimator. Numerical simulations show that the ML estimator and the proposed non-separable estimator outperform the separable Bayesian estimator for high signal-to-noise ratio (SNR). In the low SNR regime, on the other hand, the two estimators are inferior to the separable Bayesian estimator while the proposed non-separable estimator slightly outperforms the ML estimator.
@@ I-A Motivation
with noise variance $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$ for some $\sigma^{2}>0$. The signal support $\mathcal{S}_{\boldsymbol{x}}=\{n\in\mathcal{N}=\{1,\ldots,N\}:x_{n}\neq 0\}$ is assumed to be sampled from all possible sets $\mathfrak{S}_{k}^{N}=\{\mathcal{S}\subset\mathcal{N}:|\mathcal{S}|=k\}$ of size $k$ uniformly and randomly. In particular, this paper focuses on signals with sublinear sparsity: The signal sparsity kk—number of non-zero elements in $\boldsymbol{x}$—grows sublinearly in $N$ as $N\to\infty$. The goal is estimation of the sparse signal vector $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$. This simple problem arises in designing denoisers used in message-passing algorithms for sublinear sparsity [1, 2].
The noise power scaling $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$ in (1) is a key point in estimation of the sparse signal vector. Conventional research for sublinear sparsity [3, 4, 5, 6] postulated $N^{-1}\mathbb{E}[\|\boldsymbol{\omega}\|_{2}^{2}]=\Omega(1)$. This normalization implies that error-free estimation is impossible as long as the non-zero signal power is ${\cal O}(1)$. In evaluating the performance of an estimator $\hat{\boldsymbol{x}}\in\mathbb{R}^{N}$, the normalized square error $N^{-1}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}$ was used. For the scaling in this paper, on the other hand, $k^{-1}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}$ normalized by sparsity kk—simply called square error—should be considered [1, 2].
@@
When the separable Bayesian estimator is used as a denoiser, both approximate message-passing (AMP) [1] and orthogonal AMP (OAMP) [2] for sublinear sparsity require strictly larger sample complexity than information-theoretically optimal complexity [7, 8, 9, 10, 11, 12, 13]. Early research [7, 8, 9] elucidated the scaling of the optimal complexity for the linear model in compressed sensing while [10, 11] tackled generalized linear models. The prefactor in the optimal scaling was evaluated in [12, 13] for constant non-zero signals. These existing results depend strongly on the assumption of independent and identically distributed (i.i.d.) sensing matrices, so that the linear model in compressed sensing cannot be reduced to the AWGN channel (1). Thus, the existing results are not directly applicable to signal estimation from the observation in the AWGN channel (1).
The suboptimality for sublinear sparsity is in contrast to the optimality of AMP and OAMP for conventional linear sparsity $k/N\to\rho$ for some $\rho\in(0,1]$: For linear sparsity, AMP [14, 15] was proved to be Bayes-optimal [16, 17, 18] via state evolution [19, 20, 21, 22] when sparse signals are compressed with zero-mean i.i.d. sub-Gaussian matrices. Similarly, OAMP [23] or equivalently vector AMP (VAMP) [24] was proved to be Bayes-optimal [25, 26] for right-orthogonally invariant matrices via state evolution [24, 27]. See [28, 29] for a generalization of right-orthogonally invariant matrices. In contrast to these optimality for linear sparsity, both AMP and OAMP are suboptimal for sublinear sparsity [1, 2].
@@
  • In the direct part, i.e. achievability of vanishing square error, Gallager’s bound [30] is available for evaluating the performance of the maximum likelihood (ML) estimator. Gallager’s bound is a general technique that was utilized to prove achievability in sparse linear regression [11]. In contrast to a long proof for the separable Bayesian estimator [1] via a heuristic strategy, the proof of the direct part is simplified for the ML estimator.
@@
The converse part, i.e. impossibility of vanishing square error is proved via a hypothesis testing [35, 36]. In the hypothesis testing, we test whether the null hypothesis $\boldsymbol{y}\sim q(\boldsymbol{y})$ holds for some probability density function (pdf) $q(\boldsymbol{y})$ different from $p(\boldsymbol{y})$ or the alternative hypothesis $\boldsymbol{y}\sim p(\boldsymbol{y})$. When the true marginal pdf $p(\boldsymbol{y})$ of $\boldsymbol{y}$ in (1) is indistinguishable from $q(\boldsymbol{y})$ independent of the channel model (1), it is intuitively impossible to estimate the signal vector $\boldsymbol{x}$. Inspired by [13], this paper utilizes a general relationship [37] between mutual information and square error to derive a lower bound on the square error, which depends on the Kullback-Leibler (KL) divergence $D(p\|q)$ of $q(\boldsymbol{y})$ from the truth $p(\boldsymbol{y})$. In this paper, a Gaussian pdf $q(\boldsymbol{y})$ is selected. By proving that $D(p\|q)$ converges to zero, this paper proves a converse result.
@@ I-B Contributions
Before presenting theoretical results of this paper, additional assumptions are summarized. Throughout this paper, non-zero elements of the signal vector are sampled from the set of $M$ discrete non-zero points $\mathcal{U}=\{u_{m}\in\mathbb{R}\setminus\{0\}:m\in\mathcal{M}=\{1,\ldots,M\}\}$ uniformly and randomly. Let $u_{\mathrm{min}}=\min_{m\in\mathcal{M}}|u_{m}|$ and $u_{\mathrm{max}}=\max_{m\in\mathcal{M}}|u_{m}|$. The set of all possible $k$-sparse signal vectors is written as $\mathcal{X}_{k}^{N}(\mathcal{U})=\{\boldsymbol{x}\in\mathbb{R}^{N}:|\mathcal{S}_{\boldsymbol{x}}|=k,x_{n}\in\mathcal{U}~\hbox{for all $n\notin\mathcal{S}_{\boldsymbol{x}}$}\}$.
@@
Let $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ denote the ML estimator of $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$ in (1). If $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ holds, then the square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]$ converges to zero in the sublinear sparsity limit: $N$ tends to infinity with $\log k/\log N\to\gamma$ for some $\gamma\in[0,1)$.
Theorem 1 implies that the ML estimator satisfies the same achievability as the separable Bayesian estimator [1] in the low noise regime $\sigma^{2}<u_{\mathrm{min}}^{2}/2$. See Appendix A-A for a review of the separable Bayesian estimator. Since the ML estimator can be suboptimal in terms of the square error, Theorem 1 makes no claims for high noise regime. This paper presents the following converse result for high noise regime:
@@
$\gamma<\min\left\{\frac{C_{0}}{C_{0}+\sigma^{2}/(u_{\mathrm{max}}u_{\mathrm{min}})},\;\frac{C_{1}}{C_{1}+3}\right\},$ (3)
@@
We can obtain a tight result for constant non-zero signals, i.e. $u_{\mathrm{min}}=u_{\mathrm{max}}$. The upper bound in (3) is strictly positive as long as $\sigma^{2}>u_{\mathrm{min}}^{2}/2$ is satisfied. See Fig. 1 for the quantitative evaluation of $\gamma$ satisfying (3). Combining Theorems 1 and 2, this paper concludes the existence of small $\gamma$ such that the square error converges to zero in the sublinear sparsity limit if and only if $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ holds. In other words, the separable Bayesian estimator [1] is asymptotically optimal for small $\gamma$.
Figure 1: Upper bound (3) on $\gamma$ versus $u_{\mathrm{max}}^{2}/\sigma^{2}$. The lower bound (2) in Theorem 2 holds when $(u_{\mathrm{max}}^{2}/\sigma^{2},\gamma)$ is located below each curve.
The last contribution is numerical results. A non-separable estimator is proposed via a heuristic approximation of the posterior mean estimator in the noiseless limit. This estimator is called non-separable approximate Bayesian estimator or more simply non-separable estimator. The proposed non-separable estimator is numerically shown to outperform the ML estimator for all signal-to-noise ratios (SNRs) in terms of the square error. Furthermore, it is superior to the separable Bayesian estimator [1] in the high SNR regime. Unfortunately, the non-separable estimator cannot improve the performance of AMP [1] for sublinear sparsity compared to the separable Bayesian estimator. The last result suggests that AMP for sublinear sparsity has a performance bottleneck in another part rather than the type of denoisers.
@@ I-D Notation
The notation $p(X)$ denotes the pdf of an absolutely continuous random variable $X$ while $\mathbb{P}(\cdot)$ represents a probability measure. The Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ is denoted by $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. The function $Q(x)$ represents the complementary cumulative distribution function of the standard Gaussian distribution. See Table I at the end of this paper for the list of the other special notation used throughout this paper.
@@ II-A Maximum Likelihood Estimation
which minimizes the error probability $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{MAP}}\neq\boldsymbol{x})$. In terms of AMP [1] and OAMP [2] for sublinear sparsity, the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ should be used to minimize the mean square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{opt}}\|_{2}^{2}]$. Nonetheless, this paper focuses on the ML estimator (6) because the optimality of the ML estimator is partially proved in the converse part.
The ML estimator (6) can be represented via a partition of indices $\mathcal{N}=\{1,\ldots,N\}$: For some $k_{*}\in\{0,\ldots,k\}$ decompose $\mathcal{N}$ into three disjoint sets $\mathcal{N}_{1}\subset\mathcal{N}$, $\mathcal{N}_{2}\subset\mathcal{N}\setminus\mathcal{N}_{1}$, and $\mathcal{N}_{3}=\mathcal{N}\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$ with cardinality $_{*}, kk$$k_{*}$, and NkN-k, respectively. The set $\mathcal{N}_{1}$ consists of indices that correspond to the top $k_{*}$ maximum received signal values while $\mathcal{N}_{2}$ is composed of indices for the top $(k-k_{*})$ minimum received signal values. The set $\mathcal{N}_{3}$ consists of the remaining indices.
Let $\boldsymbol{y}]_{(n)}$ denote the $n$th maximum element of $\boldsymbol{y}$, i.e. $\boldsymbol{y}]_{(1)}\geq\cdots\geq[\boldsymbol{y}]_{(N)}$. Furthermore, we define the permutation $\sigma:\mathcal{N}\to\mathcal{N}$ satisfying $y_{\sigma(n)}=[\boldsymbol{y}]_{(n)}$ for all $n\in\mathcal{N}$. Using these notations, we have the representation $\mathcal{N}_{1}=\{\sigma(n):n\in\{1,\ldots,k_{*}\}\}$, $\mathcal{N}_{2}=\{\sigma(n):n\in\{N-(k-k_{*})+1,\ldots,N\}\}$, and $\mathcal{N}_{3}=\{\sigma(n):n\in\{k_{*}+1,\ldots,N-(k-k_{*})\}\}$.
@@
Determine $\mathcal{N}_{1}$, $\mathcal{N}_{2}$, and $\mathcal{N}_{3}$ with
@@ II-B Proof of Theorem 1
We evaluate the square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]$, in which the expectation is over $\boldsymbol{x}$ and $\boldsymbol{y}$ in (1). The signal vector $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$. Let $\mathcal{E}_{d}=\{\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\geq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}d\}$ with $d=\lceil k/\sqrt{\log(N/k)}\rceil$. Since $\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}$ is bounded from above by $\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}k$, we use the representation $\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]=\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}|\mathcal{E}_{d}]\mathbb{P}(\mathcal{E}_{d})+\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}|\mathcal{E}_{d}^{\mathsf{c}}]\mathbb{P}(\mathcal{E}_{d}^{\mathsf{c}})$ to obtain the following upper bound:
@@
The main challenge is evaluation of $\mathbb{P}(\mathcal{E}_{d})$. To evaluate this probability, we classify the signal space into types. Let $\boldsymbol{u}_{\boldsymbol{m}}=[u_{m_{1}},\ldots,u_{m_{k}}]^{\mathrm{T}}\in\mathcal{U}^{k}$ denote the vector of non-zero signal elements, with an index vector $\boldsymbol{m}\in\mathcal{M}^{k}$. For a subset $\mathcal{S}\subset\mathfrak{S}_{k}^{N}$ of size $k$, $\boldsymbol{m}\in\mathcal{M}^{|\mathcal{S}|}$, and $w\in\{0,\ldots,|\mathcal{S}|\}$, define the type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ of $|\mathcal{S}|$-sparse vectors in $\mathcal{X}_{|\mathcal{S}|}^{N}(\mathcal{U})$ as
@@
The type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ in (12) contains all possible $k$-sparse vectors that have non-zero elements $\boldsymbol{u}_{\boldsymbol{m}}$ and $w$ incorrectly detected positions for the signal vector $\boldsymbol{x}$. By definition, we have the following partition of the signal space $\mathcal{X}_{k}^{N}(\mathcal{U})$:
@@
We reduce the evaluation of $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$ to that of $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$ for $w^{\prime}\in\{0,\ldots,k-w\}$. The integer $w^{\prime}\in\{0,\ldots,k-w\}$ in $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ represents the number of incorrectly detected non-zero elements in correctly detected positions $n\in\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}$. Thus, this joint event implies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\leq 2u_{\mathrm{max}}^{2}w+d_{\mathrm{max}}^{2}w^{\prime}\leq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}(w+w^{\prime})$. By definition, $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$ is affected only by the events satisfying $w+w^{\prime}\geq d$ for all $w\in\{0,\ldots,k\}$ and $w^{\prime}\in\{0,\ldots,k-w\}$: